Spectrum of the linearized Euler operator

Yuri Latushkin University of Missouri-Columbia

This talk is scheduled at the Focus Session on fluid mechanics. We study spectral properties of the linearized Euler operator. For the generalized Kolmogorov flow we give the full description of the spectrum. Also, we prove spactral mapping propert that relates the spectrum of the linearized Euler operator and the spectrum of the group generated by this operator.